Calculate the percentage of premium paid given the following details:
Details:
Although rumors of the transaction leaked out yesterday, AcquirerCo officially announced today that it has agreed to buy TargetCo for $25.00 a share. TargetCo shares closed higher yesterday at $20.00.

25%

20%

15%

Not enough information

Calculate the offer price per share given the following details:
Transaction Details:
AcquirerCo agreed to buy TargetCo with a mix of cash and AcquirerCo stock. TargetCo stockholders will receive $5.00 in cash and one share of AcquirerCo common stock for every two shares of TargetCo stock. AcquirerCo's share price closed at $30.00 a day prior to the announcement.

$15.00

$35.00

$25.00

$20.00

Which form contains relevant data for an LBO in a private transaction involving non-public financing?

S-4

8-K

13E-3

None of the above

How is the equity value calculated for an M&A transaction when the target company is private?

Share price multiplied by shares outstanding

Acquirer's price per share multiplied by shares outstanding

Enterprise value less assumed/refinanced net debt

There is no equity value for a private company

What is NOT a reason why transaction comps generally provide a higher multiple range than trading comps?

Buyers pay control premiums

Buyers often have the opportunity to realize synergies

Buyers receive the right to control decisions

Transaction comps are more accurate than trading comps
Answer #1

Target company last share price =   $20
  
Price offered for Target=   $25
  
So premium offered = (New price /last price)-1  
(25/20)-1  
0.25   or 25%

(B)

(b)      
price offered to target in cash    $5  
Shares offered 1 for 2 shares held      
So Equivalent share price to Target = Acquirer share price * exchnage ratio      
30*1/2=   15  
      
So total price offered in cash and shares = 5+15=    $20  
      
      

(One separate question with parts is answered as per Chegg policy. Separate 3 Question is posted, but 1 is answered as per policy)

Ac